OpenStudy (anonymous):

A line passes through (-8,5) slope of 3/4

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

Write an equation in slope intercept form

OpenStudy (anonymous):

What do you need to know?

OpenStudy (anonymous):

Oh sorry nvm

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (jojo4eva):

y=3/4x+11

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

Thanks @jojo4eva

OpenStudy (jojo4eva):

what you do is use the equation y-y1=m(x-x1

OpenStudy (jojo4eva):

then substitute everything in and do the work and no problem